solve the following systems of linear equations using any algebraic method. If possible, check your solutions.
6. 2r - 5y = 3 Check:
x-2y = 7

Thank you for your help :)

Answer by Alan3354(69443)

Did you mean 2x ?
------
2x - 5y = 3
x - 2y = 7
---------
2x - 5y = 3
2x - 4y = 14 2nd equation times 2
------------------------------------------- Subtract
-y = -11
y = 11
---------
x - 2y = 7
x - 22 = 7
x = 29